OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately and explain each step. I.2x + 4 = 3x - 11 II. 3x + 3 = 51 III. 2x+1 all over3 = x+6 all over 2

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

@ybarrap please I don't understand

OpenStudy (ybarrap):

I. 2x + 4 = 3x - 11 2x-3x+4=3x - 11 -3x 2x + -3x+4 = - 11 -x+4 = - 11 -x+4 -4= - 11-4 -x=-15 x=15 Check: 2(15) + 4=34 3(15) - 11 = 45 - 11 = 34 II. 3x + 3 = 51 Subtract 3 from both sides and divide by 3 III. 2x+1 all over3 = x+6 all over 2 (2x + 1) / 3 = (x + 6) / 2 multiply both sides by 6 to get rid of the 3 and the 2 in the denominators (2x + 1) * 2 = (x + 6) *3 4x + 2 = 3x + 18 Subtract 2 from both sides. Subtract 3x from both sides. That's it, see which solutions match.